OpenStudy (richyw):

Show that the two-body hamiltonian\[H_{\text{sys}}=\frac{\mathbf{p}_1^2}{2m_1}+\frac{\mathbf{p}_2^2}{2m_2}+V( \mathbf{r}_1,\mathbf{r}_2)\]can be separated into centre of mass and relative hamiltonians\[H_{\text{sys}}=\frac{\mathbf{p}^2}{2M}+\frac{\mathbf{p}_{\text{rel}}}{2\mu}+V(r)\]where \(\mu\) is the reduced mass

OpenStudy (anonymous):

(I hope this is just a normal Hamiltonian and not a Hamiltonian operator) We can start out by defining the momenta in terms of mass and velocity (using the rdot nomenclature do denote dr/dt, where r is a generic position vector pointing from the origin) \[ \textbf p_1 = m_1 \dot {\textbf r}_1 \quad ; \quad \textbf p_2=m_2 \dot {\textbf r}_2\] The Center of Mass position vector is given by \[ \textbf R = \frac{m_1\textbf r_1+m_2\textbf r_2}{M}\] And taking the first derivative wrt time gives the velocity of the CoM as \[ \dot{\textbf R}=\frac{m_1\dot{\textbf r}_1+m_2\dot{\textbf r}_2}{M}\] And similarly the momentum of the CoM as \[\textbf{P}_{CoM}=\textbf P = M\dot{\textbf R}\] Next, the relative position vector r is defined to be \[ \textbf r = \textbf r_1-\textbf r_2\] Next you can rearrange the CoM position vector and the relative position vectors with a bit of algebra (about 4 lines) to give \[ \textbf r_1 = \textbf R + \frac{m_2}{M} \textbf r \quad \quad \text{and} \quad \quad \textbf r_2 = \textbf R - \frac{m_1}{M} \textbf r\] and the first derivative wrt time is \[ \dot{\textbf r}_1 = \dot{\textbf R} + \frac{m_2}{M}\dot{\textbf r} \quad \quad \text{and} \quad \quad \dot{\textbf r}_2 = \dot{\textbf R} - \frac{m_1}{M}\dot{\textbf r}\] Looking again at our Hamiltonian, we can first say that it is a function of r1dot and r2dot and r1 and r2 \[H_{(\dot{ \textbf r}_1 , \dot{ \textbf r}_2, \textbf r_1 , \textbf r_2)}=T_{(\dot{ \textbf r}_1 , \dot{ \textbf r}_2)}+V_{( \textbf r_1,\textbf r_2)}\] Looking at the kinetic energy portion \[T_{(\dot{ \textbf r}_1 , \dot{ \textbf r}_2)}=\frac{1}{2}\Big( m_1 \dot{ \textbf r}_1^2 + m_2 \dot{ \textbf r}_2^2 \Big)\] Now we can substitute in our values for r1dot and r2dot in terms of the first derivative of the relative position terms \[T_{(\dot{\textbf R}, \dot{\textbf r})} = \frac{1}{2} \left( m_1 \left[ \dot{ \textbf R} + \frac{m_2}{M} \dot{ \textbf r} \right]^2+ m_2 \left[\dot{ \textbf R} - \frac{m_1}{M} \dot{ \textbf r} \right]^2 \right) \] \[=\frac{1}{2}\left((m_1+m_2)\dot{ \textbf R}^2+(m_1+m_2)\frac{m_1 m_2}{M^2}\dot{ \textbf r}^2 \right)\] etc. Reduce it a few more times and that gives the kinetic energy in terms of the CoM and the relative momenta. Remember that \[ M = m_1+m_2\] \[\mu = \frac{m_1m_2}{M}\] \[ \textbf P = M\dot{ \textbf R}\] \[ \textbf P_{rel} = \mu \dot{ \textbf r}\] \[T_{(\dot{\textbf R},\dot{\textbf r})}=T_{(\dot{\textbf R})}+T_{(\dot{\textbf r})}\] As for the potential V, since our position vectors r1 and r2 are pointing from the origin (a requirement of the CoM equation) then the distance between 2 objects we're looking at - be it masses or charges - is merely the magnitude of the vector pointing from one object to the other - aka, the magnitude of our relative position vector \[\textbf r = \textbf r_1-\textbf r_2\] so for example \[ V_{(\textbf r_1 , \textbf r_2)} = -G\frac{m_1m_2}{|r_1-r_2|}\] gives \[V_{(\textbf r_1 , \textbf r_2)} = V_{(|\textbf r_1 - \textbf r_2|)}\] aka \[V=V_{(r)}\] So now we can separate the Hamiltonian into the CoM and relative components \[ H = H_{CoM} + H_{rel}\] \[ = T_{\dot{(\textbf R})} + (T_{(\dot{\textbf r})}+V_{(r)})\] Just finish reducing the kinetic energy, and you're good to go ^_^